# End-to-End Predictive Planner for Autonomous Driving with Consistency Models

### v1

@@ V Methodology
With the aforementioned notations and stated problem, our proposed method is summarized in Fig. 1. We first use an MTR encoder to encode the agent’s trajectory history and map information. Then we use a consistency model that takes the conditional input of the ego agent’s planning goal and the MTR-encoded features and generates trajectory plans and predictions for the ego and surrounding agents, respectively. Additional planning constraints for the ego agent are achieved through guided sampling of consistency models.
@@ V-A Data pre-processing
For training diffusion models or consistency models, we usually need to normalize the input data. If we jointly model ego and surrounding agents’ future trajectories $(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})$, using an ego-centric coordinate system introduces large variances in the data, particularly for surrounding agents whose positions vary significantly across scenarios. Thus the training performance is largely degraded. Instead, we apply a coordinate transformation $\Gamma$ similar to MTR++ [46] that maps each agent’s trajectory into its own local coordinate frame, centered at its position at the current timestep $t_{0}$. This transformation provides us with data $(\Gamma(S_{e}^{\text{future}}),\Gamma(S_{\mathcal{O}_{s}}^{\text{future}}))\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times H_{2}\times d_{s}}$ that has substantially reduced variance. We then compute the empirical mean and standard deviation of the transformed trajectories across the dataset and standardize them to zero mean and unit variance. To preserve the relative spatial relationships between agents, we collect the reference states consisting of each agent’s position at timestep $t_{0}$ as $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times d_{s}}$.
@@ V-B Consistency Model Training
Let $\mathcal{X}$ denote the space of transformed future trajectories of the ego vehicle $\Gamma(S_{e}^{\text{future}})$ and surrounding agents $\Gamma(S_{\mathcal{O}s}^{\text{future}})$ from Sec. V-A. Let $\mathcal{Y}$ denote the space of conditional inputs, containing encoded historical trajectories $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})$, map features $\text{MTR}(M)$ from Sec. IV-A, ego agent’s goal states $s_{e}^{g}$, and the reference coordinates $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})$ from Sec. V-A.
Given a planning goal and environment context as the condition $y\in\mathcal{Y}$, we leverage the consistency model [52, 50] to sample future trajectories from the conditional probability distribution $p(\cdot|y)$ over $\mathcal{X}$. Each trajectory sample represents one possible joint future behavior of the ego vehicle and surrounding agents.
For training our predictive planner, we jointly train the MTR encoder and the consistency model together in one step with a hybrid loss function, which is constructed as a weighted sum of the consistency training loss $\mathcal{L}_{\text{consistency}}$ in Eq. (IV-B) and the dense prediction loss $\mathcal{L}_{\text{encoder}}$ of the MTR encoder from Sec. IV-A as follows:
@@
V-C Guided Sampling
When generating trajectories with the trained consistency model as described in Sec. IV-B, to impose planning constraints on the ego vehicle’s future trajectory, we present a novel guided sampling approach similar to classifier guidance [13]. Importantly, this guidance is applied solely during sampling at test time, without requiring any modification to the training procedure.
Assuming there are $N_{c}\in\mathbb{N}$ planning constraint $c_{j}$ to minimize, $j=1,2,...,N_{c}$ and drawing inspiration from classifier guidance [13], we can perform a gradient descent step on the predicted $x_{1}^{[i]}$ during each sampling step $i$ in Eq. (4):
$x_{1}^{[i,1]}=x_{1}^{[i]}-\sum^{N_{c}}_{j=1}\alpha_{j}\nabla c_{j}$ (6)
However, optimizing multiple constraints simultaneously presents a significant challenge in finding a suitable stepsize $\alpha$ efficiently, particularly when constraints may conflict with each other. To address this challenge, inspired by the Alternating Direction Method of Multipliers (ADMM) [6], we propose a novel alternating direction method that only optimizes one constraint at a time during each sampling step $i$:
$x_{1}^{[i,1]}=x_{1}^{[i]}-\alpha_{1}\nabla c_{1}$
$x_{1}^{[i,2]}=x_{1}^{[i,1]}-\alpha_{2}\nabla c_{2}$
$x_{1}^{[i,N_{c}]}=x_{1}^{[i,N_{c}-1]}-\alpha_{N_{c}}\nabla c_{N_{c}}$ (7)
In our approach, we sequentially optimize each constraint $c_{j}$ with corresponding step sizes $\alpha_{j}$ during each gradient descent iteration. While our selection of the optimization order and relatively small step sizes is based on empirical observations, this strategy has demonstrated effective convergence in practice, which is difficult to achieve with the standard guidance method with gradient descent in Eq. (6). The convergence of alternating direction methods, such as ADMM, is generally guaranteed under conditions including closed, proper, and convex functions and appropriately chosen step sizes [6].
V-D Planning Constraints Construction
We assume that the ego vehicle’s dynamics are as follows:
@@ V-D Planning Constraints Construction
where $p_{x}$ and $p_{y}$ are positions, $v$ is the linear speed and $\theta$ is the orientation. The control input is the acceleration $a$ and angular speed $\omega$.
Suppose we only use $p_{x}$ and $p_{y}$ from ego agent’s future states $S_{e}^{\text{future}}$ to construct our planning constraint function $c_{i}$, since other states like $v$ or $\theta$ may not satisfy the dynamics equations in Eq. (8) and are also quite noisy. Using the property of differential flatness for dynamics in Eq. (8), we can infer $v$ and $\theta$ with $p_{x}$ and $p_{y}$ as follows:
$v=\sqrt{\dot{p}_{x}^{2}+\dot{p}_{y}^{2}},\quad\theta=\arctan(\frac{\dot{p}_{y}}{\dot{p}_{x}})$ (9)
Then we can further infer the control input $a$ and $\omega$:
$a=\frac{\dot{p}_{x}\ddot{p}_{x}+\dot{p}_{y}\ddot{p}_{y}}{\sqrt{\dot{p}_{x}^{2}+\dot{p}_{y}^{2}}},\quad\omega=\frac{\dot{p}_{x}\ddot{p}_{y}-\dot{p}_{y}\ddot{p}_{x}}{\dot{p}_{x}^{2}+\dot{p}_{y}^{2}}$ (10)
Note that, $a$ and $v$ are computed through finite differencing of $p_{x}$ and $p_{y}$ which will inevitably incur some noises.
With $p_{x},p_{y}$ and the derived $a$ and $\omega$, we consider three types of planning constraint $c$ to minimize as follows.
@@
$c_{\text{goal}}$ represents a goal-reaching constraint such that the ego vehicle needs to reach a pre-specified goal $(p_{x,\text{goal}},p_{y,\text{goal}})$ at the final timestep $t_{0}+H_{2}$. $c_{\text{acc}}$ and $c_{\omega}$ are control limit constraints that compute the average control limit violation along the ego agent’s trajectory for a given $a_{\text{limit}}$ and $\omega_{\text{limit}}$.
@@ VI-A1 Dataset
We evaluate our method on the Waymo Open Motion Dataset (WOMD) [15], which contains 486k traffic scenarios in the training set and 44k scenarios in the validation set that we can test on. Each scenario provides road graph information and trajectories of diverse agents including vehicles, pedestrians, and cyclists, with up to 8 agents specifically marked for trajectory prediction. For each agent, this dataset includes 1 second of history, current state, and 8 seconds of future trajectory, sampled at 0.1-second intervals. This sets our history horizon to be $H_{1}=10$ and future horizons to be $H_{2}=80$ timesteps, respectively. For each scenario, with the trajectory information from all agents including vehicles, pedestrians, and cyclists as $S_{\mathcal{O}}$, we only select vehicle agents as the ego agent with trajectories denoted by $S_{e}$. We then identify the surrounding agents, with trajectories denoted as $S_{\mathcal{O}_{s}}$, by selecting up to 4 agents closest to the ego vehicle based on their trajectory distances, excluding those beyond a 10-meter threshold. This enables us to focus our predictive planner on the most relevant interactive behaviors between the ego vehicle and its immediate surroundings. The goal $s_{e}^{g}$ of the ego agent is chosen as its state at the end of the 8-second future horizon.
@@ VI-A2 Model Details
The MTR encoder consists of 6 transformer layers with a dimension of 256 to encode agent trajectory histories and map information into $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}},M)$. For detailed encoder architecture specifications, we refer readers to [45]. The conditional input $y$ for the consistency model is constructed by further encoding three components: the obtained features $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}},M)$, the ego agent’s goal $s_{e}^{g}$, and the reference states $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})$, using 3-layer MLPs with 256 neurons per layer. To model the consistency model function $F_{\theta}$, we adopt the U-Net [40] with the model dimension of 128 as the backbone.
For the consistency model, we choose $T=5$ which indicates 4 sampling steps in the forward and reverse process. Follow the original paper [50], we set the noise schedule with $\rho=6$, $\sigma_{1}=0.002,\sigma_{T}=80$, and then $\sigma_{i}=\left(\sigma_{0}^{1/\rho}+\frac{i-1}{T-1}(\sigma_{T}^{1/\rho}-\sigma_{0}^{1/\rho})\right)^{\rho},i=1,...,T$. During guided sampling, we perform 100 gradient steps of Eq. (V-C) at each sampling step, and set the coefficient for $c_{\text{goal}},c_{\text{acc}}$ and $c_{\omega}$ to be $2\times 10^{-5},3\times 10^{-6},5\times 10^{-7}$, respectively.
Figure 2: Ego vehicle’s planned trajectory (Green) in a left-turn scenario compared to the Groundtruth (Red).
TABLE I: Planning Trajectory Compared to the Human Groundtruth
Method minADE $\downarrow$ minFDE $\downarrow$
Transformer 0.944 3.239
DDPM-10 0.346 0.377
DDPM-4 0.363 0.498
DDIM-4 2.126 2.559
Consistency 0.309 0.218
Consistency (guided) 0.303 0.016
(a) A proactive nudging behavior from the ego car to switch a lane.
(b) A yielding behavior for the ego car to make a turn.
@@ VI-A3 Methods and Baselines
We compare our consistency model-based approach against several baselines. Currently, we only provide the evaluation of the open-loop trajectory generation performance of those methods, focusing on the behavior through a single inference step.
@@
A model based on MTR’s encoder-decoder architecture [45] to generate trajectory only for ego vehicle. It has a larger model size for the decoder (512 vs. our 128), but doesn’t incorporate ego vehicle’s goal $s_{e}^{g}$ as input.
DDPM
This diffusion baseline uses the same MTR encoder and MLP architecture as ours for condition encoding but adopts DDPM [20] for trajectory sampling of ego and surrounding agents. Note that it differs from consistency models in both its forward diffusion process and training loss function. We evaluate two variants: one trained and sampled with 4 steps (DDPM-4), and another trained and sampled with 10 steps (DDPM-10).
@@
This baseline extends DDPM-10 by incorporating the more efficient DDIM sampling strategy [48]. While using the same training process as DDPM-10, it employs 4-step DDIM sampling during inference, denoted as DDIM-4.
Consistency
Our proposed consistency model is trained and sampled with 4 steps.
Consistency (guided)
Our proposed consistency model is trained and sampled with 4 steps, with guided sampling.
@@ VI-A4 Training
We conduct distributed training with a total batch size of 100 across 4 NVIDIA L40 GPUs, each with a batch size of 25. For consistency and diffusion models, we use the Adam optimizer [29] with a learning rate of $8\times 10^{-5}$ and train for 50 epochs. For the transformer, we follow the MTR settings [45] and use the AdamW optimizer [35] with a learning rate of $1\times 10^{-4}$ and weight decay of 0.01, and train for 30 epochs.
@@ VI-B Experiment results and analysis
In this section, we evaluate our consistency model-based planner from several aspects. First, we assess its planning performance and trajectory quality compared to several baseline methods. Then we take a closer look at interactive scenarios and present examples of proactive nudging and yielding behaviors of our predictive planner. Furthermore, we show how online guided sampling improves constraint satisfaction in planning - the controllability not provided with transformer-based approaches. Finally, we discuss the computational time during inference for our presented methods.
VI-B1 Trajectory Planning Performance
In Fig. 2, we compare planned trajectories generated by different methods (shown in green) against the ground truth from the dataset (shown in red). The scenario requires the ego vehicle to make a left turn and then drive straight to reach its goal. According to Fig. 2, DDPM-4 reaches the goal but produces a noisy trajectory, likely due to insufficient sampling steps for diffusion models to generate high-quality trajectories. While DDPM-10’s increased sampling steps yields slightly smoother trajectories, it comes at the cost of longer computation time. DDIM-4 attempts to accelerate DDPM-10 using only 4 sampling steps, but fails to generate goal-reaching trajectories. The transformer with a larger model size generates a trajectory close to the groundtruth but misses the exact goal location. In contrast, our consistency model generates smooth trajectories that both align with the ground truth and precisely reach the target location.
@@ VI-B1 Trajectory Planning Performance
In Table I, we quantitatively evaluate the planning performance for the ego vehicle compared to the human groundtruth through metrics from Waymo’s Motion Prediction Challenge [58]: Minimum Average Displacement Error (minADE) and Minimum Final Displacement Error (minFDE) on the entire test dataset. Here we view these metrics as an evaluation of how closely our generated trajectories align with human driving behavior under the same environmental context and goal conditions. The results align with our qualitative observations: our consistency model achieves the lowest minADE and minFDE, indicating that it captures human driving patterns well. Moreover, the addition of guided sampling significantly improves the minFDE for the consistency models since it explicitly incorporates goal-reaching constraints.
@@
In Table II, we evaluate the trajectory quality using three metrics that characterize driving behavior: 1) angle changes, which measure the smoothness of directional transitions, 2) path length, which quantifies the efficiency of the chosen route, and 3) curvature, which assesses the sharpness of turns. Lower values in these metrics generally indicate smoother and more efficient trajectories that align with natural driving patterns. Our consistency model achieves superior trajectory quality compared to all diffusion-based methods, including DDPM-10 despite its additional sampling steps. while the transformer method has slightly lower angle changes and curvature, this is expected given its substantially larger model size (512 vs. our 128 dimensions).
TABLE II: Planning Trajectory Quality
Method Angle changes $\downarrow$ Path len $\downarrow$ Curvature $\downarrow$
Transformer 0.419 59.621 0.560
DDPM-10 0.660 87.801 3.444
DDPM-4 0.729 79.731 2.211
DDIM-4 0.683 157.78 2.012
Consistency 0.636 58.060 1.174
Consistency (guided) 0.445 57.564 0.610
@@
Furthermore, we demonstrate the flexibility of our approach in choosing goal location in Fig. 3. Our consistency model maintains high-quality trajectory planning when targeting novel goal locations, even when they are not present in the dataset. The current transformer method with MTR doesn’t support the selection of goal locations.
VI-B2 Interactive Behavior
A key advantage of our predictive planner is its ability to ensure safe and effective interactions with other road users in traffic. Our planner demonstrates the capability for generating both proactive behaviors as shown in Fig. 4, such as nudging for lane switching and yielding to pedestrians. The ego agent (labeled “E”) trajectory is in green and the surrounding agents (labeled “S”) trajectories are in orange. Other agent’s trajectories are shown in blue. These learned behaviors are similar to ground truth human drivers and ensure efficient navigation and safe interaction between the ego vehicle and other road users. Crucially, all planning and prediction are achieved through a single inference of the consistency model.
In Fig. 4a, we show a proactive nudging scenario. Here, the ego vehicle needs to merge into the left lane while a nearby vehicle occupies that lane. Our planner directs the ego vehicle to accelerate and nudge into the left lane, predicting that the vehicle in the left lane will decelerate to accommodate the lane change. This nudging behavior demonstrates the proactive planning capability of our end-to-end, data-driven predictive planner, which is challenging to achieve using modular methods. While in the ground truth, the left car doesn’t decelerate that much and the rightmost vehicle eventually turns right, this discrepancy is expected given that we currently generate 8-second future trajectories through a single inference.
In Fig. 4b, a yielding behavior is presented. The ego vehicle intends to make a right turn but predicts that a pedestrian will continue crossing the road. As a result, the planner instructs the ego vehicle to wait until the pedestrian has fully crossed, which is highly aligned with the ground truth. This yielding behavior ensures safe interaction between the ego vehicle and other road users, highlighting the planner’s safe operation in dynamic environments.
VI-B3 Guided Sampling
The guided sampling approach described in Sec. V-C and Sec. V-D allows us to enforce planning constraints during inference time, providing a key advantage over transformer-based methods which lack this capability. We quantitatively evaluate all methods against three planning constraints from Eq. (V-D): goal reaching, acceleration limits, and orientation speed limits. As shown in Table III, our base consistency model already achieves more accurate goal-reaching and smoother behaviors than other diffusion-based approaches. With guided sampling, these planning constraint violations of the consistency model further decrease by a factor of 4, achieving the lowest values among all methods. Note that this improvement requires no model retraining and only applies during inference.
TABLE III: Planning Constraint Violations
Method Goal Reaching $\downarrow$ Acc limit $\downarrow$ $\omega$ limit $\downarrow$
Transformer 9.256 0.570 1.029
DDPM-10 12.669 80.581 3.307
DDPM-4 13.873 56.170 3.323
DDIM-4 22.933 218.823 5.270
Consistency 0.618 5.754 2.050
Consistency (guided) 0.127 0.462 0.599
VI-B4 Computational time discussions
@@
All methods in our evaluation support batch sampling. As shown in Table IV, we evaluate their inference times using a batch of 30 scenarios with 125 ego agents on an NVIDIA GeForce RTX 3060. While the transformer achieves the fastest inference time, it lacks our method’s key capabilities: online controllable generation and sampling from multimodal joint distribution. Our consistency model achieves a similar inference time with DDPM-4 and DDIM-4 due to the same model dimensionality, and sampling steps. However, DDPM-4 and DDIM-4 fail to match our consistency model’s performance in trajectory generation across all metrics, as evidenced in Fig. 2, Table I, Table II, and Table III. The guided sampling, introducing a little computational overhead, provides higher quality trajectory generation and enables controllable planning. Even DDPM-10, despite its increased sampling steps and higher computational time, cannot achieve a comparable performance of trajectory generation to our approach.
TABLE IV: Computational Time Comparison for Inference
Method Inference Time (s)
Transformer 1.910
DDPM-10 3.352
DDPM-4 2.226
DDIM-4 2.118
Consistency 2.301
Consistency (Guided) 2.635
VII Limitations
Although our proposed method significantly improves the efficiency of generating human-like trajectories through consistency models and joint prediction and planning, it may not yet meet the demands of real-time operation, especially in highly dynamic environments that require high-frequency planning. Future work could leverage model pruning or distillation techniques to further improve the inference efficiency. Besides, while in the current paper, we test open-loop trajectory generation to focus on evaluating the model performance of a single inference step, in the future we will extend our method to the closed-loop case with a receding horizon fashion, which is expected to further enhance the interactive planning.
VIII Conclusions
In this paper, we present a novel data-driven approach that unifies trajectory prediction and planning using consistency models. By modeling the joint distribution of ego and surrounding agents’ trajectories, our method enables efficient and interactive planning, such as proactive nudging and yielding, without the need for iterative alternation between prediction and planning modules. Trained on the Waymo Open Motion Dataset, the proposed consistency model generates human-like trajectories while its guided sampling capability naturally incorporates additional planning constraints during online trajectory generations. Experimental results demonstrate that our approach achieves superior trajectory quality and interactive behavior compared to existing methods while requiring fewer sampling steps. This framework also has the potential to be extended to broader robot decision-making tasks in interactive scenarios.
Looking forward, we aim to optimize this method for real-time operation and extend it to closed-loop operation with receding horizon planning. Additionally, we plan to incorporate LLMs as high-level planners to interpret diverse contextual information and enhance decision-making in challenging scenarios.

### v2

@@ V Methodology
Our method is summarized in Fig. 1. We first use an MTR encoder to encode the agent’s trajectory history and map information. With the ego agent’s planning goal and the MTR-encoded features, a consistency model generates trajectory plans and predictions for the ego and surrounding agents, respectively. Multiple planning constraints for the ego agent are achieved through ADMM-inspired online guided sampling.
@@ V-A Data pre-processing
We apply a coordinate transformation $\Gamma$ similar to MTR++ [35] that maps each agent’s trajectory into its own local coordinate frame, centered at its position at the current timestep $t_{0}$. The transformed data $(\Gamma(S_{e}^{\text{future}}),\Gamma(S_{\mathcal{O}_{s}}^{\text{future}}))\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times H_{2}\times d_{s}}$ has substantially reduced variance. We then standardize the transformed trajectories to zero mean and unit variance. To preserve the relative spatial relationships among agents, the reference states are collected as $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})\in\mathbb{R}^{(N_{\mathcal{O}_{s}}+1)\times d_{s}}$ consisting of each agent’s position at timestep $t_{0}$.
@@ V-B Consistency Model Training
(a) A proactive nudging behavior from the ego car to switch a lane.
(b) A safe yielding behavior for the ego car to make a turn.
Let $\mathcal{X}$ denote the space of transformed future trajectories $\Gamma(S_{e}^{\text{future}})$ and $\Gamma(S_{\mathcal{O}s}^{\text{future}})$ from Sec. V-A. Let $\mathcal{Y}$ denote the space of conditional inputs, containing encoded historical trajectories $\text{MTR}(S_{e}^{\text{hist}},S_{\mathcal{O}}^{\text{hist}})$, map features $\text{MTR}(M)$ from Sec. IV-A, ego agent’s planning goal $s_{e}^{g}$, and the reference coordinates $\text{Ref}(S_{e}^{\text{future}},S_{\mathcal{O}_{s}}^{\text{future}})$ from Sec. V-A. Given a sample $y\in\mathcal{Y}$, we leverage the consistency model [40, 39] to sample future trajectories from the conditional probability distribution $p(\cdot|y)$ over $\mathcal{X}$. Each trajectory sample represents one possible joint future behavior of the ego vehicle and surrounding agents.
For training our predictive planner, we jointly train the MTR encoder and the consistency model together with a hybrid loss function, which is a weighted sum of the consistency training loss $\mathcal{L}_{\text{consistency}}$ in Eq. (IV-B) and the dense prediction loss $\mathcal{L}_{\text{encoder}}$ of the MTR encoder from Sec. IV-A as follows:
@@
V-C Planning Constraints Construction
To make the sampled trajectories suitable and smooth for planning use, we apply several constraints on the ego trajectories. Assuming the vehicle has the following dynamics
@@ V-C Planning Constraints Construction
where $p_{x}$ and $p_{y}$ are positions, $v$ is the linear speed and $\theta$ is the orientation. The control input is the acceleration $a$ and angular speed $\omega$. We only use $p_{x}$ and $p_{y}$ from the future states $S_{e}^{\text{future}}$ to construct our planning constraint function $c_{i}$, since other states like $v$ or $\theta$ in the dataset may not satisfy the dynamics equations in Eq. (4) and are also quite noisy. Using the differential flatness for dynamics in Eq. (4), we infer the control input $a$ and $\omega$ as $a=\frac{\dot{p}_{x}\ddot{p}_{x}+\dot{p}_{y}\ddot{p}_{y}}{\sqrt{\dot{p}_{x}^{2}+\dot{p}_{y}^{2}}},\omega=\frac{\dot{p}_{x}\ddot{p}_{y}-\dot{p}_{y}\ddot{p}_{x}}{\dot{p}_{x}^{2}+\dot{p}_{y}^{2}}$ through finite differencing with inevitable noises.
With $p_{x},p_{y}$ and the derived $a$ and $\omega$, we consider three types of planning constraints $c$ to minimize as follows.
@@
$c_{\text{goal}}$ enforces goal-reaching for the ego vehicle at the final timestep $t_{0}+H_{2}$. $c_{\text{acc}}$ and $c_{\omega}$ are control limit constraints over the ego agent’s trajectory for a given $a_{\text{limit}}$ and $\omega_{\text{limit}}$. These constraints are inherently in conflict. For example, only moving the trajectory’s final state closer to the goal can increase acceleration violations of the car
V-D Guided Sampling
To simultaneously enforce multiple planning constraints, like in Eq. (V-C), on the trajectory samples without affecting the original distribution, we present a novel ADMM-inspired guided sampling approach, which is applied solely during sampling at test time, without requiring any modification to the training procedure.
Assuming there are $N_{c}\in\mathbb{N}$ planning constraint $c_{j}$ to minimize, $j=1,2,...,N_{c}$, standard classifier-guidance [11] performs a vanilla gradient descent step on the predicted $x_{1}^{[i]}$ during each sampling step $i$ in Eq. (2) as follows:
$x_{1}^{[i,1]}=x_{1}^{[i]}-\sum^{N_{c}}_{j=1}\alpha_{j}\nabla c_{j}$ (6)
However, optimizing multiple constraints simultaneously with conflicts presents a challenge in finding a suitable stepsize $\alpha$. To address this, we propose a novel alternating direction method inspired by ADMM [6] that sequentially applies gradient step for each constraint during the sampling step $i$:
$x_{1}^{[i,j+1]}=x_{1}^{[i,j]}-\alpha_{j}\nabla c_{j},\quad\text{for}\quad j=1,...,N_{c}$ (7)
We find that this alternating update strategy demonstrates effective convergence and empirically leads to lower violations of multiple conflicting constraints, compared to standard guided sampling with vanilla gradient descent in Eq. (6). It also simplifies step size tuning.
@@ VI-A Experiment setup
TABLE I: Planning Trajectory Realism
Method minADE ($m$) minFDE ($m$) Step
Transformer 1.237 2.796 1
Diffusion-10 0.488 0.532 10
Diffusion-4 0.545 0.711 4
DDIM 2.749 2.304 4
Ours 0.497 0.173 4
Ours (guided) 0.490 0.008 4
TABLE II: Planning Trajectory Quality and Constraint Violations for Different Methods
Method Planning Trajectory Quality Planning Constraint Violations
Angle changes Path length Curvature Collision rate Goal reaching Acceleration limit $\omega$ limit
[…]
Transformer 0.465 55.885 0.599 15.0 8.111 0.438 1.119
Diffusion-10 0.639 67.899 1.392 1.7 2.410 35.609 2.248
Diffusion-4 0.716 61.407 1.379 2.0 3.171 19.812 2.566
DDIM 0.673 135.55 1.755 18.6 14.440 175.602 4.545
Ours 0.670 55.373 1.107 0.8 0.361 4.806 2.064
Ours (guided) 0.487 55.137 0.556 0.8 0.026 0.281 0.519
TABLE III: Increased constraint violation (%) for vanilla gradient descent compared to the ADMM-inspired method.
Gradient Step Goal reaching Acceleration limit $\omega$ limit
50 10% 26% 8%
100 9% 35% 8%
150 12% 41% 9%
@@ VI-A1 Dataset
We evaluate our method on the Waymo Open Motion Dataset (WOMD) [13], which contains 486k traffic scenarios in the training set and 42k scenarios in the interactive validation set to test on. For each agent, including vehicles, pedestrians, and cyclists, the dataset includes 1 second of history, current state, and 8 seconds of future trajectory, sampled at 0.1-second intervals. This sets our history horizon to be $H_{1}=10$ and future horizons to be $H_{2}=80$ timesteps. We identify the surrounding agents with trajectories $S_{\mathcal{O}_{s}}$ by selecting up to 4 agents closest to the ego vehicle based on their trajectory distances, excluding those beyond a 10-meter threshold. This enables us to focus our predictive planner on the most relevant interactive behaviors between the ego vehicle and its immediate surroundings. The goal $s_{e}^{g}$ of the ego agent is chosen as its state at the end of the 8-second future horizon.
@@ VI-A2 Model Details
The MTR encoder consists of 6 transformer layers with a dimension of 256 to encode agent trajectory histories and map information [34]. The consistency model’s conditional embedding uses 3-layer MLPs with 256 neurons per layer. The consistency function $F_{\theta}$ is modeled with a U-Net [30] with a dimension of 128.
We choose $T=5$, i.e., 4 sampling steps in the forward and backward process for the consistency model. We set the noise schedule with $\rho=6$, $\sigma_{1}=0.002,\sigma_{T}=80$, and $\sigma_{i}=\left(\sigma_{0}^{1/\rho}+\frac{i-1}{T-1}(\sigma_{T}^{1/\rho}-\sigma_{0}^{1/\rho})\right)^{\rho},i=1,...,T$ as  [39]. During guided sampling, we perform 100 gradient steps of Eq. (7) at each sampling step, and set the coefficient for $c_{\text{goal}},c_{\text{acc}}$ and $c_{\omega}$ to be $2\times 10^{-5},3\times 10^{-6},5\times 10^{-7}$, respectively.
@@ VI-A3 Methods and Baselines
We compare our consistency model-based approach against several baselines. Currently, we provide the evaluation of the open-loop trajectory generation performance of those methods, focusing on interactive behavior through a single inference.
@@
This transformer baseline uses MTR [34] to generate trajectory only for ego vehicle. It has a larger model size for the decoder (512 vs. our 128), but doesn’t incorporate the ego vehicle’s goal $s_{e}^{g}$ as input.
Diffusion-4 and Diffusion-10
These diffusion baselines use the same MTR encoder and goal embedding as our approach but adopt DDPM [16] for trajectory sampling. We evaluate two variants: one trained and sampled with 4 steps (Diffusion-4), and another with 10 steps (Diffusion-10).
@@
This baseline aims to accelerate the pretrained Diffusion-10 model with only 4 sampling steps in trajectory generation, with an efficient DDIM sampling strategy [37].
Ours and Ours guided
These are our proposed consistency model trained and sampled with 4 steps, with and without the proposed guided sampling.
@@ VI-A4 Training
We conduct distributed training with a total batch size of 100 across 4 NVIDIA L40 GPUs. For consistency and diffusion models, we use the Adam optimizer [24] with a learning rate of $8\times 10^{-5}$ and train for 50 epochs. For the transformer, we follow the MTR settings [34] and use the AdamW optimizer [24] with a learning rate of $1\times 10^{-4}$ and weight decay of 0.01, and train for 30 epochs.
@@ VI-B Experiment results and analysis
VI-B1 Interactive Behavior
We demonstrate that our predictive planner can generate safe and effective interactions with other road users with a single inference of the consistency model, without alternation between prediction and planning.
In Fig. 2, we show a proactive nudging and a safe yielding behavior of our planner, with the ego agent (“E”) trajectory in green, the surrounding agents (“S”) in orange, and other agents in blue. In Fig. 2a, our planner directs the ego vehicle to accelerate and nudge into the left lane, predicting that the vehicle in the left lane will decelerate to accommodate the lane change. While in the ground truth, the rightmost vehicle eventually turns right instead of going straight, this discrepancy is expected for an 8-second future prediction. In Fig. 2b, the planner instructs the ego vehicle to wait until the pedestrian has fully crossed before making the right turn.
VI-B2 Planning Performance under Limited Computational Budget
Figure 4: Constraint violation losses with respect to gradient steps during the final sampling step of the consistency model. Vanilla gradient descent struggles with three conflicting constraints, while the ADMM-inspired method optimizes smoothly.
@@ VI-B2 Planning Performance under Limited Computational Budget
In Table I, we quantitatively evaluate the realism of the planned trajectory for the ego vehicle through metrics with human groundtruth: Minimum Average Displacement Error (minADE) and Minimum Final Displacement Error (minFDE) [45] over 6 sampled trajectories. The results show that our consistency model, with only 4 sampling steps, captures realistic human driving patterns comparable to Diffusion-10. Moreover, the addition of guided sampling significantly improves the minFDE for the consistency models since it explicitly incorporates goal-reaching constraints.
@@
In Table II, we evaluate the trajectory quality using four metrics 1) angle change as the smoothness of directional transitions, 2) path length as the efficiency of the chosen route, 3) curvature as the sharpness of turns, and 4) collision rate ($2m$ threshold) as safety. Lower values in these metrics indicate smoother, safer, and more efficient trajectories that align with natural driving patterns. Our consistency model achieves superior trajectory quality compared to all Diffusion-based methods, including Diffusion-10, despite its additional sampling steps. While the transformer method has slightly lower angle changes, this is expected given its substantially larger model size (512 vs. our 128 dimensions).
@@
Furthermore, we demonstrate the flexibility of our approach in choosing goal location in Fig. 3. Our consistency model maintains high-quality trajectory planning when targeting novel goal locations not present in the dataset. The current transformer method (MTR) doesn’t support goal selections.
@@
With the same encoder, the computational time for diffusion-based methods directly depends on the number of sampling steps, shown in Table I. Diffusion-4 and DDIM have the same sampling step as our model, but fail to match our performance in trajectory generation across all metrics, as evidenced in Table I and Table II. The guided sampling, introducing a little computational overhead ($\approx 15\%$), provides higher quality trajectory generation and enables controllable planning. Even Diffusion-10, despite its more sampling steps, cannot achieve a comparable performance of trajectory generation to our approach in Table II.
VI-B3 Ablation Study on ADMM-inspired Guided Sampling
We compare our ADMM-inspired method (ADMM for short) in Eq. (7) with the vanilla gradient descent (GD for short) in Eq. (6), which simultaneously optimizes all planning constraints. To illustrate the advantage of our method in hyperparameter tuning, we sweep through all combinations of step sizes ($\alpha=1e^{-4},1e^{-5},1e^{-6}$) across three constraints (27 combinations total), using 50, 100, and 150 gradient steps over 200 scenarios. TABLE III shows that GD results in significantly higher violations than ADMM on average; e.g., at 100 steps, GD has 9% higher goal-reaching, 35% higher acceleration, and 8% higher angular speed violations. The lowest sums of violations achieved are 1.45, 1.16, and 0.71 for 50, 100, and 150 steps, respectively. Increasing gradient steps reduces constraint violations but with computational tradeoffs.
In Fig. 4, we visualize how the violation loss evolves with ADMM (blue) and GD (red) methods in the last guided sampling step, measured across 200 scenarios and with the same step size finally chosen. As discussed in Sec. V-D, these three constraints are in conflict with each other. In Fig. 4, GD experiences large oscillations in acceleration violations (middle plot), highlighting its struggles with optimizing conflicting objectives. In contrast, our ADMM method smoothly and effectively reduces all three violations.
VII Conclusion
We propose a consistency model-based predictive planner that generates safe and highly interactive behavior, such as nudging and yielding, for autonomous vehicles. Compared to diffusion models and their accelerated variants like DDIM, our method generates smoother, safer, and more efficient paths under a limited computational budget. To satisfy multiple planning constraints online, we introduce a novel ADMM-inspired approach used in guided sampling for consistency models. Future work will explore the integration of LLMs as high-level planners for closed-loop decision-making in long-tailed challenging scenarios.